# GP-GS: Gaussian Processes for Enhanced Gaussian Splatting

### v3

@@ A.1 MOGP Training and Dynamic Filtering Procedure
Input: Image size $(W,H)$, depth map $D$, training pixels $\{(u_{i},v_{i})\}_{i=1}^{N}$, angular resolution $M$, scale factor $\beta$, variance quantile threshold $R^{2}$.
@@
    Compute adaptive sampling radius $r=\beta\cdot\min(H,W)$.
@@
    for each training pixel $(u_{i},v_{i})$:
    • Generate $M$ new samples in a circular neighborhood.
    • Normalize $(u,v)$ and retrieve depth $d$ from $\mathcal{D}$.
    • Store $\bigl{(}u,v,d\bigr{)}$ in sample set $\mathcal{\tilde{P}}$.
@@
Step 4: Final Point Cloud $\mathcal{\tilde{P}}^{*}\leftarrow\mathcal{\tilde{P}}\cup\widehat{\mathcal{P}}$.
@@
A.2 MOGP Training Loss
In this section, we present the training loss curves of our MOGP model across different datasets. The loss curves illustrate the convergence behaviour of the model during training, providing insight into its stability and optimization progress. Figure 7 shows the loss curves for various datasets over 1,000 iterations, including bicycle, garden, room, lego, kitchen, truck, flower, and drums. Our MOGP model demonstrates a smooth and consistent decline in loss, indicating effective learning and convergence.
A.3 More MOGP Results
Table 5: Performance Metrics for MOGP Model Using Different Matern Kernel Parameters. The best results are highlighted. One can see that $\nu=0.5$ helps to produce the best densification results across various datasets.
@@ A.3 More MOGP Results
In this section, we present additional experimental results to further evaluate the effectiveness of our MOGP model. Specifically, we analyze the impact of different Matérn kernel parameters on performance across various datasets and assess the effectiveness of our variance-based filtering approach. To validate our uncertainty-based filtering algorithm, we examine the variance distributions before and after applying the filter.
Evaluation of $\nu$ values on Pixel-to-Point-Cloud dataset. The Matérn kernel introduces a smoothness parameter $\nu$, enabling adjustable local variations depending on the scene. Through iterative gradient updates, we refine the kernel hyperparameters to enhance both accuracy and generalization. We show the details on various datasets in Table 5 to choose the best performance $\nu$ value for the kernel.
@@
Effectiveness of Variance-Based Filtering. Figure 8 illustrates that our filter effectively removes high-variance noisy points while maintaining scene consistency. Additional results on various datasets, including NeRF Synthetic (Lego), Tanks & Temples (Truck), and Mip-NeRF 360 (Flowers), are presented. The right column displays the variances after applying our filter. The variance in spatial components $(x,y,z)$ and colour channels $(r,g,b)$ is significantly reduced, demonstrating the filter’s efficacy in removing high-variance noisy points while preserving scene consistency.
Figure 8: Variance reduction comparison of our variance-based filtering Algorithm. The left and right columns display the variances before and after applying our filter. One can see that the variance in spatial components $(x,y,z)$ and RGB channels $(r,g,b)$ is significantly reduced, demonstrating the filter’s effectiveness in eliminating high-variance noise while preserving scene consistency and structural integrity.
These results combined with the results we presented in this paper further demonstrate the robustness and generalizability of our proposed MOGP-based point cloud densification method. By leveraging uncertainty estimation, our approach ensures a more structured and reliable reconstruction of complex 3D scenes.
A.4 More Novel View Synthesis Results
Figure 9: Further Qualitative comparison examples of ours GP-GS and 3DGS*. One can see that GP-GS reconstructs intricate structures like branches, trunks, and leaves more accurately, while 3DGS* produces blurry results. Similar results are observed in the truck and kitchen scenes, i.e., GP-GS better preserves fine details, reducing artefacts and improving spatial consistency.
In this section, we present more qualitative comparison results of GP-GS and 3DGS*.
@@ A.4 More Novel View Synthesis Results
As show in Figure 9. Our GP-GS (right) outperforms 3DGS* (middle) in complex regions with densely packed objects (e.g., foliage) and challenging lighting conditions. Notably, in the flower scene, GP-GS reconstructs intricate structures like branches, trunks, and leaves more accurately, while 3DGS* produces blurry results. Similarly, in the truck and kitchen scenes, our method better preserves fine details, reducing artefacts and improving spatial consistency. These results further validate the effectiveness of GP-GS in enhancing 3D Gaussian splitting for reconstruction.

### v4

@@ IV-C Point Cloud Densification
Input: Image size $(W,H)$, depth map $\mathcal{D}$, training pixel coordinates $\{[u_{i},v_{i}]^{T}\}_{i=1}^{N}$, angular resolution $M$, scale factor $\beta$, variance quantile threshold $R^{2}$.
@@
    Compute adaptive sampling radius $r=\beta\cdot\min(W,H)$.
@@
    for each training pixel with coordinates $[u_{i},v_{i}]^{T}$:
    • Generate $M$ new samples stacked as $[\mathbf{u},\mathbf{v}]^{T}\in\mathbb{R}^{M\times 2}$ in a circular neighborhood.
    • Normalize $[\mathbf{u},\mathbf{v}]^{T}$ and retrieve depth $\mathbf{d}\in\mathbb{R}^{M}$ from $\mathcal{D}$.
    • Store $\bigl{[}\mathbf{u},\mathbf{v},\mathbf{d}\bigr{]}^{T}$ in sample set $\mathcal{\tilde{P}}$.
@@
Step 4: Final Point Cloud $\mathcal{\tilde{P}}^{*}\leftarrow\mathcal{P}\cup\widehat{\mathcal{P}}$.
@@
V Experiments
V-A Dataset and Implementation Details
Dataset. Previous approaches to SfM often overlook the explicit correspondence between 2D image pixels and sparse 3D point clouds, we address this gap by introducing the Pixel-to-Point-Cloud dataset, which will be opensourced along the codes. This dataset provides a detailed mapping between image pixels and their corresponding 3D point clouds, thus facilitating more accurate modelling and enabling the training of our MOGP model. We further conduct a comprehensive evaluation of our method on dataset spanning NeRF Synthetic [19], Mip-NeRF 360 [1], and Tanks and Temples [14]. These benchmarks are popular in 3D reconstruction and novel-view synthesis due to their diverse characteristics, including high-fidelity 3D models, large-scale outdoor scenes with intricate scenes like clustered leaves, and lighting condition variations. By leveraging our new dataset and evaluation framework, we demonstrate the effectiveness of our approach in advancing point cloud densification and improving reconstruction fidelity across varied and challenging scenarios.
Implementation Details The MOGP-based densification process in GP-GS can be regarded as a plug-and-play module and can be integrated into any 3D reconstruction framework based on SfM sparse point clouds. We have integrated the process into 3DGS*, one of the state-of-the-art 3D reconstruction models, and compared their performance over the datasets mentioned earlier following 3DGS*’s training hyperparameters.
Our MOGP models are trained for 1,000 iterations across all scenes, we set L2 regularization weight $\lambda=10^{-6}$, learning rate $l=0.01$, dynamic sampling resolution $M=8$, adaptive sampling radius $r=0.25$. All experiments are conducted on an RTX 4080 GPU.
Metrics. The evaluation of the GP-GS is twofold, i.e., the evaluation of the SfM sparse point densification, and the evaluation of the novel view rendering performance. For densification, in line with prior studies on point cloud reconstruction and completion [37, 39], we adopt the mean Chamfer Distance (CD) to measure the discrepancy between the predicted point clouds and the ground truth. Specifically, for a predicted point set $\mathcal{P}$ and its corresponding ground truth set $\mathcal{G}$, the CD between them is calculated as:
We also compute the Root Mean Squared Error (RMSE) and $R^{2}$ [7] to show the robustness of GP-GS under various metrics. The RMSE is calculated to quantify the average squared deviation between the predicted and true values, using $\mathrm{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(\mathbf{g_{i}}-\mathbf{p_{i}})^{2}}$, a lower RMSE value indicates a more accurate prediction. The $R^{2}$ score captures the proportionate reduction in the residual variance:
For novel view rendering performance evaluation, we compare our method against state-of-the-art NVS approaches based on commonly used image-based metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [30], and Learned Perceptual Image Patch Similarity (LPIPS) [41] on the rendered images in the test views.
V-B Results of Multi Output Gaussian Process
Quantitative Results
Figure 3: Variance reduction comparison of our variance-based filtering Algorithm. The left and right columns display the variances before and after applying our filter. We can see that the variance in spatial components $(x,y,z)$ and RGB channels $(r,g,b)$ is significantly reduced, demonstrating the filter’s effectiveness in eliminating high-variance noise while preserving scene consistency and structural integrity.
To ensure the best performance of MOGP for densification, we evaluated MOGP model with different hyperparameters on our Pixel-to-Point-Cloud dataset, e.g., $\nu$ with values such as ($\nu\in\{0.5,1.5,2.5\}$). The data were split into an 80% training subset and a 20% testing subset, and the model was subsequently assessed using $R^{2}$, RMSE, and CD as we mentioned in section V-A. The average outcomes of this analysis are summarized in Table I, which shows that $\nu=0.5$ consistently yields the most favourable results. Consequently, we adopt $\nu=0.5$ as our default configuration for subsequent experiments. Detailed metrics across multiple datasets for this kernel selection procedure are presented in Table II.
TABLE I: Comparison of different Matern kernel smoothness parameters $\nu$. The best results are highlighted. best $R^{2}$, best RMSE, and best CD.
$\nu$ Value $R^{2}\uparrow$ $\mathrm{RMSE}\downarrow$ $\mathrm{CD}\downarrow$
0.5 0.71 0.13 0.21
1.5 0.67 0.15 0.24
2.5 0.61 0.14 0.26
TABLE II: Performance Metrics for MOGP Model Using Different Matern Kernel Parameters. The best results are highlighted. best $R^{2}$, best RMSE, and best CD. We can see that $\nu=0.5$ helps to produce the best densification results across various datasets.
@@ V-B Results of Multi Output Gaussian Process
To validate the effectiveness of our variance-based filtering algorithm, we analyze the variance distributions across multiple Pixel-to-Point datasets, including NeRF Synthetic, Tanks & Temples, and Mip-NeRF 360, due to their diverse characteristics. The reduction in variance in all datasets suggests that our filtering method generalizes well to different types of 3D scenes, improving overall point cloud quality and enabling more accurate scene representation, as shown in Figure 3, the left column shows the spatial and colour variances before filtering on NeRF Synthetic (lego), Mip NeRF 360 (flowers) and Tanks & Temples (Truck), the right column presents the variances after applying our filter. The variances in spatial components $(x,y,z)$ and colour channels $(r,g,b)$ are significantly reduced, demonstrating that our filtering process effectively removes high-variance noisy points while preserving scene consistency.
MOGP Training Loss We also present the training loss curves of our MOGP model across different datasets. The loss curves illustrate the convergence behaviour of the model during training, providing insight into its stability and optimization progress. Figure 4 shows the loss curves for various datasets over 1,000 iterations, including bicycle, garden, room, lego, kitchen, truck, flower, and drums. Our MOGP model demonstrates a smooth and consistent decline in loss, indicating effective learning and convergence.
Figure 5: Qualitative results comparison of our MOGP point cloud densification approach and SfM.
Figure 6: Qualitative comparison of 3D Gaussians during 3DGS training process. The left column shows the ground truth images. The middle column presents 3D Gaussians from 3DGS*, while the right column shows 3D Gaussians generated using ours MOGP.
Further validating our densified point cloud, we present Figure 6, which shows 3D Gaussians during training. Our approach leads to more structured and coherent reconstructions, especially in complex regions with fine textures and occlusions. In the flower scene, our 3D Gaussian representation successfully reconstructs branches, trunks, and leaves on the left and right sides, while the original 3D Gaussian sphere remains blurry and lacks structural detail. Additionally, our method accurately preserves the flower shape in the middle, improving overall reconstruction fidelity. In the room scene, our method provides a clearer representation of the curtain folds on the left, whereas the original method produces irregularly shaped Gaussian spheres and visible artefacts. The highlighted insets further demonstrate that our method reduces visual artefacts and better preserves geometric consistency during training.
V-C Quantitative and Qualitative Results on NVS
Figure 7: Qualitative comparisons of novel view synthesis on three benchmark datasets: Tanks & Temples (Truck), Mip-NeRF 360 (flowers and bicycles), and NeRF Synthetic (lego). From left to right: Ground Truth, 3DGS*, Ours (MOGP + 3DGS*), Scaffold-GS, and Ours (MOGP + Scaffold-GS). This Figure highlight key differences in texture sharpness, structural coherence, and occlusion boundaries. Our method consistently recovers finer details (e.g., leaves, spokes, and edges) and preserves global geometry better across all scenes. These results further validate the effectiveness of our proposed plug-and-play MOGP module, which improves both 3DGS* and Scaffold-GS.
To evaluate the effectiveness and versatility of our MOGP method, we conduct comprehensive experiments on datasets Tanks & Temples, Mip-NeRF 360, and NeRF Synthetic. Our proposed MOGP model is designed to be plug-and-play, enabling seamless integration into existing novel view synthesis pipelines. We incorporate MOGP into both 3DGS* [13] and Scaffold-GS [18], and compare performance using both quantitative metrics and qualitative visualizations.
Quantitative Results Table III,  IV,  V report PSNR, SSIM, and LPIPS scores across all datasets, respectively. Our method consistently achieves higher PSNR and SSIM values while maintaining lower LPIPS scores, indicating superior reconstruction quality and perceptual realism. The improvements are particularly evident in regions with high structural complexity and varying textures, such as dense foliage and intricate surfaces, where existing methods often suffer from over-smoothing and aliasing artefacts. Our approach effectively preserves fine details and enhances geometric consistency, even under challenging lighting conditions. Figure 7 provides a detailed visual comparison.
TABLE III: Comparison of Various Methods on Tanks and Temples. Best results are highlighted: best PSNR, best SSIM, and lowest LPIPS.
Dataset Scaffold-GS Ours (Scaffold-GS + MOGP) 3DGS* Ours (3DGS* + MOGP)
[…]
Truck 25.45 0.879 0.150 25.57 0.880 0.150 25.10 0.876 0.153 25.18 0.877 0.153
Train 22.50 0.828 0.198 22.53 0.830 0.196 22.13 0.819 0.202 22.22 0.823 0.199
Average 23.98 0.854 0.174 24.05 0.856 0.173 23.62 0.847 0.178 23.70 0.850 0.176
TABLE IV: Comparison of Various Methods on NeRF Synthetic. Best results are highlighted: best PSNR, best SSIM, and lowest LPIPS.
Dataset Scaffold-GS Ours (Scaffold-GS + MOGP) 3DGS* Ours (3DGS* + MOGP)
[…]
chair 30.89 0.976 0.027 30.92 0.976 0.027 30.99 0.978 0.024 30.99 0.978 0.024
drums 25.51 0.936 0.068 25.53 0.937 0.068 25.52 0.940 0.065 25.52 0.940 0.065
ficus 25.98 0.925 0.061 25.99 0.927 0.060 25.03 0.923 0.058 25.20 0.923 0.058
hotdog 34.51 0.978 0.042 34.61 0.978 0.041 34.53 0.979 0.037 34.55 0.979 0.037
lego 28.62 0.926 0.075 28.90 0.929 0.074 28.32 0.929 0.074 29.08 0.936 0.067
materials 17.88 0.749 0.221 18.11 0.755 0.219 16.82 0.734 0.239 16.68 0.736 0.241
mic 27.67 0.852 0.153 27.71 0.853 0.152 18.90 0.855 0.144 19.02 0.863 0.123
ship 27.66 0.853 0.153 27.65 0.852 0.151 27.08 0.848 0.157 27.05 0.849 0.154
Average 27.34 0.899 0.100 27.43 0.901 0.099 25.90 0.898 0.100 26.01 0.901 0.096
TABLE V: Comparison of Various Methods on Mip-NeRF 360. Best results are highlighted: best PSNR, best SSIM, and lowest LPIPS.
Dataset Scaffold-GS Ours (Scaffold-GS + MOGP) 3DGS* Ours (3DGS* + MOGP)
[…]
bicycle 19.70 0.643 0.262 19.80 0.644 0.260 19.17 0.632 0.248 19.22 0.634 0.248
bonsai 18.25 0.588 0.445 18.50 0.593 0.437 18.66 0.579 0.458 18.69 0.581 0.455
counter 29.73 0.917 0.121 29.83 0.918 0.119 29.15 0.911 0.122 29.17 0.915 0.121
garden 29.56 0.919 0.064 29.53 0.919 0.064 29.10 0.915 0.063 29.38 0.920 0.061
kitchen 31.86 0.943 0.073 31.87 0.944 0.073 31.74 0.943 0.069 32.28 0.948 0.068
room 32.22 0.939 0.118 32.30 0.939 0.120 31.78 0.934 0.126 31.89 0.939 0.121
flowers 23.08 0.707 0.237 23.15 0.709 0.236 23.04 0.707 0.251 23.16 0.711 0.246
Average 26.34 0.808 0.189 26.43 0.809 0.187 26.09 0.803 0.191 26.26 0.807 0.189
@@ V-C Quantitative and Qualitative Results on NVS
Qualitative Results As shown in Figure 7, we compare our method with 3DGS* on datasets Tanks & Temples (Truck), Mip-NeRF 360 (Flowers and Bicycles), and NeRF Synthetic (Lego). Our GP-GS (right) outperforms 3DGS* (middle) in complex regions with densely packed objects (e.g., foliage) and challenging lighting conditions. Notably, in the flower scene, GP-GS reconstructs intricate structures like branches, trunks, and leaves more accurately, while 3DGS* produces blurry results. Similarly, in the truck and bicycle scenes, our method better preserves fine details, reducing artifacts and improving spatial consistency. These results further validate the effectiveness of GP-GS in enhancing 3D Gaussian splatting for reconstruction. Our method demonstrates superior performance in NVS by effectively reconstructing fine-grained details, reducing artifacts, and enhancing occlusion handling. The visual comparisons highlight key challenges in 3DGS* rendering, including occlusions, high-frequency textures (e.g., object edges and surface details), and lighting variations. Notably, the zoom-in regions (marked with red and blue boxes) illustrate our model’s ability to recover fine structures more accurately than existing approaches. These results further validate the robustness of our method in handling complex scenes, offering improved perceptual quality and structural consistency across diverse datasets.
VI Conclusion
We Propose GP-GS, a novel framework that enhances 3DGS and improves rendering quality by addressing the limitations of sparse SfM reconstructions. Our approach formulates point cloud densification as a continuous regression task using a Multi-Output Gaussian Process (MOGP), which adaptively densifies sparse SfM point clouds by learning mappings from 2D image pixels and depth priors to 3D positions and colors, guided by predictive uncertainty. Furthermore, we introduce an adaptive neighborhood-based sampling strategy to improve point cloud completeness while ensuring structured densification. To further refine the densified point clouds, we incorporate a variance-based filtering method, effectively removing high-uncertainty predictions and preserving geometric consistency. Extensive experiments on synthetic and real-world datasets show that GP-GS significantly improves 3D reconstruction quality, and can provide compact 3D Gaussian representation, especially in complex regions with densely packed objects or under challenging lighting conditions. Future work includes integrating temporal information for dynamic scenes and scaling to larger environments.